### v2

@@ 3.5 Iterative Latent Refinement
where temperature $\tau=\min(0.5,e^{0.001j})$ decreases with refinement step $j$, gradually making the latent representation aware of quantization. Notably, selective compression is not incorporated during SGA. Instead, the selective compression map is predicted by $f_{s}$ based on the overfitted latent representations obtained after SGA. we primarily employ it as a validation tool to demonstrate our model’s potential for instance-level adaptability and performance improvements. Rate control mechanisms [84] could be incorporated in future work for applications requiring strict bit-rate constraints.
@@
   Kodak    Tecnick    CLIC Pro Valid
   PSNR    MS-SSIM    PSNR    PSNR
   Cheng’20 [16]    CVPR’20    5.58    –44.21    7.57    11.71
   ChARM [51]    ICIP’20    3.23    –    –0.88    –
   Qian’21 [55]    ICLR’21    10.05    –39.53    7.52    0.02
   Xie’21 [75]    ACMMM’21    1.55    –43.39    -0.80    3.21
   NLAIC’21 [15]    TIP’21    11.80    –40.27    7.11    –
   Gao’21 [24]    ICCV’21    –8.76    –48.60    -    –6.49
   Guo’22 [26]    TCSVT’22    –4.45    –45.23    –    –
   LBHIC [74]    TCSVT’22    –4.56    –50.54    –    –
   Entroformer [54]    ICLR’22    4.73    –42.64    2.31    –1.04
   SwinT-ChARM [87]    ICLR’22    –1.73    –    –    –
   NeuralSyntax [71]    CVPR’22    8.97    –39.56    –    5.64
   STF [88]    CVPR’22    –2.48    –47.72    –2.75    0.42
   WACNN [88]    CVPR’22    –2.95    –47.71    –5.09    0.04
   ELIC [30]    CVPR’22    –5.95    –44.60    –9.14    –3.45
   Contextformer [38]    ECCV’22    –5.77    –46.12    –9.05    –
   GLLMM [22]    TIP’23    –1.43    –47.41    –5.58    –
   NVTC [21]    CVPR’23    –1.04    –    –    –
   LIC-TCM Large [45]    CVPR’23    –10.14    –48.94    –11.47    –8.04
   QARV [19]    TPAMI’24    0.31    –    –3.03    -
   FTIC [40]    ICLR’24    –12.99    –51.13    –14.88    –9.53
   LLIC-TCM [32]    TMM’24    –10.94    –49.73    –14.99    –10.41
   WeConvene [23]    ECCV’24    –6.71    –48.63    –8.42    –5.74
   CCA [27]    NeurIPS’24    –12.04    –    –14.55    –10.75
   MambaIC [82]    CVPR’25    –13.03    –    –18.41    –15.58
   LALIC [20]    CVPR’25    –13.72    –50.86    –18.19    –14.34
   DCAE [47]    CVPR’25    –15.43    –53.20    –20.51    –15.78
   Baseline Models
   MLIC [34]    ACMMM’23    -8.05    –49.13    –12.73    –8.79
   MLIC+ [34]    ACMMM’23    -11.39    –52.75    –16.38    –12.56
   MLIC++ [33]    ICMLW’23    -13.39    -53.63    -17.59    -13.08
   MLICv2    Ours    -16.54    -54.56    -21.61    -16.05
   MLICv2+    Ours    -20.46    -55.86    -24.35    -19.14
Figure 7: Reconstructions of our proposed MLICv2, recent learned image compression models [33, 30, 88, 16] and VTM-17.0 Intra [12]. “[MSE]” denotes the model is optimized for MSE. Please zoom in for better view.
(1) In MLICv2+∗, gradient checkpointing [14] is employed during the SGA iterations.
(2) The MACs of MLICv2+ are theoretically estimated based on the number of SGA iterations, following the observation from Epoch AI that backward and forward passes almost always have MAC ratios close to 2:122footnotemark: 2. The MACs of MLICv2+∗ are not reported due to the use of gradient checkpointing, which prevents accurate profiling.
Figure 8: GPU memory consumption, Forward MACs, encoding time and decoding time comparisons among our proposed models and recent learned image compression models  [16, 75, 71, 88, 30, 45, 40, 34, 33].
@@ 4.1.1 Training Dataset
Our training dataset contains 101,104 images with resolutions exceeding $640\times 640$ pixels. These images are carefully selected from various established image datasets [18, 43, 1, 42, 77]. To address the existing compression artifacts in JPEG images, we follow the approach of Ballé et al. [10] by further down-sampling the JPEG images using a randomized factor. Additionally, images with bits per pixel (bpp) lower than 3 are excluded to ensure data quality.
@@ 4.1.2 Training Strategy
MLICv2 is implemented using PyTorch 2.2.2 and CompressAI 1.2.6. Following CompressAI, we configure $\lambda\in\{18,35,67,130,250,483\}\times 10^{-4}$ for MSE optimization and $\lambda\in\{2.4,4.58,8.73,16.64,31.73,60.5\}$ for Multi-Scale Structural Similarity (MS-SSIM) [73]. Training is conducted on eight Tesla A100-80G GPUs with a batch size of 16, employing the Adam optimizer $\beta_{1}=0.9,\beta_{2}=0.999$. Our training consists of two stages: Stage 1: We train the model with only hyperprior for 2 million (M) steps using a learning rate of $10^{-4}$ and $256\times 256$ patches, providing superior initialization for analysis and synthesis transforms. Stage 2: Loading the pre-trained weights, we train the complete model for 2M steps. The learning rate follows a scheduled decay: starting at $10^{-4}$, then decreasing to $3\times 10^{-5}$, $10^{-5}$, $3\times 10^{-6}$ and $10^{-6}$, at 1.5M, 1.8M, 1.9M, and 1.95M steps, respectively. We employ $256\times 256$ patches for the initial 1.2M steps, then transition to $512\times 512$ patches to leverage the effectiveness of global context modules. For the guided selective compression module optimization, we conduct 20K steps with batch size 16 and learning rate $10^{-3}$.
@@ 4.2 Benchmarks and Metrics
We evaluate rate-distortion performance on three datasets extensively used in learned compression research [9, 64, 10, 50, 51, 16, 75, 31, 30, 88, 87, 38, 19, 71, 26, 74, 15, 45, 34, 33, 32]:
  • Kodak [37]: Contains 24 uncompressed images with $768\times 512$ pixels.
  • Tecnick [4]: Comprises 100 uncompressed images with $1200\times 1200$ pixels.
  • CLIC Pro Valid [65]: The validation set from the 3rd Challenge on Learned Image Compression, consisting of 41 high-resolution images with $2048\times 1440$ pixels.
Performance ranking employs the Bjøntegaard delta rate (BD-Rate) metric [11] for comprehensive evaluation.
@@ 4.3.1 Quantitive Results
For a fair comparison with learned image compression models, VTM-17.0 Intra is evaluated in YUV444 color space under encoder_intra_vtm.cfg. To demonstrate the superiority of our MLICv2 and MLICv2+, we compare with recent learned image compression models [41, 16, 74, 51, 55, 75, 15, 26, 54, 87, 88, 30, 38, 22, 21, 45, 40, 32, 23, 20, 47].
The rate-distortion curves are presented in Fig. 6. When compared with our baseline MLIC++, MLICv2+ performs significantly better, especially at high bit-rates and on high-resolution images [4, 65]. Specifically, our MLICv2 achieves a maximum improvement of $0.6,0.5,0.5$ dB when the bpp is around $0.8,0.55,0.6$ on Kodak, Tecnick, CLIC Pro Val, respectively. Our MLICv2+ achieves larger improvement compared to MLIC++ [33] due to adopted latent refinement [80] for enhanced instance adaptability. The average improvements of latent refinement on MLICv2 on Kodak [37], Tecnick [4], CLIC Pro Val [65] are $0.22,0.15,0.15$ dB, respectively.
BD-Rate reductions over VTM-17.0 Intra are presented in Table 1. Compared with existing methods, our MLICv2 and MLICv2+ achieve state-of-the-art performance, reducing BD-Rates by $16.54\%$ and $20.46\%$ on Kodak. When compared to the baseline model MLIC++ [33], our MLICv2 and MLICv2+ reduce $3.15\%$ and $7.07\%$ more bit-rates over VTM-17.0 Intra on Kodak [37], respectively. Our MLICv2 and MLICv2+ also achieve superior performances on high-resolution datasets Tecnick [4] and CLIC Pro Val [65]. Specifically, our MLICv2 and MLICv2+ reduce BD-Rates by $21.61\%$ and $24.35\%$ on Tecnick [4] and $16.05\%$ and $19.14\%$ on CLIC Pro Val [65].
@@ 4.3.2 Qualitative Results
We also compare the qualitative performance to MLIC++, ELIC [30], STF, WACNN [88], Cheng’20 [16] and VTM-17.0 Intra on subjective qualities. The windowsill of reconstructions are cropped to patches for clearer comparisons. The reconstructions of MLICv2 have sharper textures and retain more details. In terms of visual quality, our MLICv2 has significant improvements on rate-perception performance compared to other models.
4.4 Computational Complexity
The model complexities and sizes are presented in Fig. 8 and Table 5, respectively. To comprehensively evaluate model complexity, 16 images from the LIU4K dataset [44] are cropped to $\{512\times 512,768\times 768,1024\times 1024,1536\times 1536,2048\times 2048,2560\times 2560,3072\times 3072\}$ patches to cover the various resolutions of images that can be encountered in practical scenarios. Forward GPU memory, encoding time, decoding time, and forward Multiply-Accumulate operations (MACs) are employed as the complexity metrics. Encoding and decoding times include arithmetic encoding and decoding times. Since the number of skipped elements at different bit-rates varies, we report the results of MLICv2 and MLICv2+ with $\lambda=0.0130$.
Due to the high memory consumption introduced by SGA, we employ gradient checkpointing [14] to overcome GPU memory limitations. This technique alleviates memory pressure by selectively recomputing a subset of intermediate activations during the backward pass. Instead of storing all intermediate activations during the forward pass, only a small set—typically those essential for gradient computation—are cached. The remaining activations are recomputed on-the-fly during backpropagation. This strategy significantly reduces peak memory usage, albeit at the cost of increased computation time.
@@ 4.4 Computational Complexity
11footnotetext: https://epoch.ai/blog/backward-forward-FLOP-ratio
@@ 4.4.1 On Forward GPU Memory
Compared to MLIC+[34], MLICv2 significantly reduces memory usage thanks to its linear-complexity entropy model. Relative to MLIC++[33], it incurs slightly higher memory for high-resolution images due to the additional modules introduced in this work, but still requires less memory than Xie’21 [75], Qian’21 [55], Entroformer [55], and LIC-TCM [45]. This modest increase is justified by the corresponding performance improvements. Latent refinement, while increasing memory during backpropagation, is optional and primarily benefits scenarios where encoding complexity is less critical. Since images are typically encoded once but decoded multiple times, the added encoding cost is acceptable given the notable gains in rate-distortion performance.
@@
Transform Architecture Cheng’20 [16] SwinT [87] ELIC [30] MLIC++ LIC-TCM [45] MLICv2
BD-Rate(%) 0.00 -4.61 -5.19 -0.41 -8.91 -9.44
Table 2: Comparisons among different transforms of proposed MLICv2, LIC-TCM [45], MLIC++, ELIC [30], and Cheng’20 [16]. All method are quiped with autoregressive context module [52].
@@
Context Module MLIC++ MLICv2 Hyperprior Guided Global Context 2D RoPE Context Reweighting Selective Compression
@@
4.4.2 On Forward MACs
Compared to MLIC+, LIC-TCM [45], our MLICv2 has lower forward MACs. Compared to MLIC++, our MLICv2 exhibits slightly higher forward MACs. This increase can be attributed to two primary factors: the proposed channel reweighting module and selective compression module. However, considering the significant performance gains achieved through these enhancements, the marginal increase in MACs is deemed acceptable. The trade-off between computational complexity and improved compression efficiency is well-balanced in our proposed model.
@@ 4.4.3 On Encoding Time and Decoding Time
Compared to MLIC and MLIC+ [34], MLICv2 achieves higher encoding and decoding efficiency due to its linear-complexity components. Relative to MLIC++[33], it further reduces encoding and decoding time by efficiently skipping many zero elements, which is particularly beneficial for high-resolution images where zeros are more frequent. Latent refinement, while involving multiple backpropagation steps and increasing encoding time, is optional and suited for scenarios where encoding complexity is less critical. Since images are typically encoded once but decoded multiple times, the extra encoding cost is justified by the performance gains. Notably, MLICv2+ maintains decoding complexity nearly identical to MLICv2, with both substantially faster than MLIC++, recent WeConvene [23] and DCAE [47].
@@ 4.4.4 On Model Size
Context Modules MLICv2+ Cumulative Benefits Individual Contributions
Full Case 1 Case 2 Case 3 Case 4 Case 5 Case 6 Case 7 Case 8 Case 9 Case 10 Case 11 Case 12 Case 13 Case 14 Case 15 Case 16
Latent Refinement
Selective Compression
@@
Context Reweighting
Hyperprior Guided Global
Simple Token Mixing Transform
@@
BD-Rate (%) 0.00 0.17 4.89 5.21 5.77 7.79 8.29 9.58 9.87 4.64 9.52 9.12 7.64 8.77 8.51 9.58 9.87
Table 4: Complete ablation study on Kodak dataset. BD-Rate (%) is employed to evaluate component contributions. MLICv2+ is the anchor. Cases 1-8 show cumulative benefits, Cases 9-16 show individual contributions.

### v3

@@ 3.5 Iterative Latent Refinement (ILR)
where temperature $\tau=\min(0.5,e^{0.001j})$ decreases with refinement step $j$, gradually making the latent representation aware of quantization. Notably, selective compression is not incorporated during SGA. Instead, the selective compression map is predicted by $f_{s}$ based on the overfitted latent representations obtained after SGA. we primarily employ it as a validation tool to demonstrate our model’s potential for instance-level adaptability and performance improvements. Rate control mechanisms zhang2024practical could be incorporated in future work for applications requiring strict bit-rate constraints.
@@
Kodak kodak Tecnick tecnick2014TESTIMAGES CLIC Pro Valid CLIC2020
PSNR MS-SSIM PSNR MS-SSIM PSNR MS-SSIM
MLIC jiang2022mlic ACMMM’23 -8.05 -49.13 -12.73 -47.26 -8.79 -45.79
MLIC+ jiang2022mlic ACMMM’23 -11.39 -52.75 -16.38 -53.54 -12.56 -48.75
MLIC++ jiang2023mlicpp ICMLW’23 -13.39 -53.63 -17.59 -53.83 -13.08 -50.78
MLICv2 Ours -16.54 -54.56 -21.61 -55.26 -16.05 -51.86
MLICv2+ Ours -20.46 -55.86 -24.35 -56.42 -19.14 -52.75
@@ 4.1.1 Training Dataset
Our training dataset contains 101,104 images with resolutions exceeding $640\times 640$ pixels. These images are carefully selected from various established image datasets deng2009imagenet, lin2014microsoft, Agustsson2017NTIRE2C, lim2017enhanced, yang2023hq. To address the existing compression artifacts in JPEG images, we follow the approach of Ballé et al. balle2018variational by further down-sampling the JPEG images using a randomized factor. Additionally, images with bits per pixel (bpp) lower than 3 are excluded to ensure data quality.
@@ 4.1.2 Training Strategy
MLICv2 is implemented using PyTorch 2.2.2 and CompressAI 1.2.6. Following CompressAI, we configure $\lambda\in\{18,35,67,130,250,483\}\times 10^{-4}$ for MSE optimization and $\lambda\in\{2.4,4.58,8.73,16.64,31.73,60.5\}$ for Multi-Scale Structural Similarity (MS-SSIM) wang2003multiscale. Training is conducted on eight Tesla A100-80G GPUs with a batch size of 16, employing the Adam optimizer $\beta_{1}=0.9,\beta_{2}=0.999$. Our training consists of two stages: Stage 1: We train the model with only hyperprior for 2 million (M) steps using a learning rate of $10^{-4}$ and $256\times 256$ patches, providing superior initialization for analysis and synthesis transforms. Stage 2: Loading the pre-trained weights, we train the complete model for 2M steps. The learning rate follows a scheduled decay: starting at $10^{-4}$, then decreasing to $3\times 10^{-5}$, $10^{-5}$, $3\times 10^{-6}$ and $10^{-6}$, at 1.5M, 1.8M, 1.9M, and 1.95M steps, respectively. We employ $256\times 256$ patches for the initial 1.2M steps, then transition to $512\times 512$ patches to leverage the effectiveness of global context modules. For the guided selective compression module optimization, we conduct 20K steps with batch size 16 and learning rate $10^{-3}$.
@@ 4.2 Benchmarks and Metrics
We evaluate rate-distortion performance on three datasets extensively used in learned compression research balle2016end, theis2017lossy, balle2018variational, minnen2018joint, minnen2020channel, cheng2020learned, xie2021enhanced, he2021checkerboard, he2022elic, zou2022the, zhu2022transformerbased, koyuncu2022contextformer, duan2023qarv, wang2022neural, guo2021causal, wu22block, chen2021nic, liu2023learned, jiang2022mlic, jiang2023mlicpp, jiang2024llic:
  • Kodak kodak: Contains 24 uncompressed images with $768\times 512$ pixels.
  • Tecnick tecnick2014TESTIMAGES: Comprises 100 uncompressed images with $1200\times 1200$ pixels.
  • CLIC Pro Valid CLIC2020: The validation set from the 3rd Challenge on Learned Image Compression, consisting of 41 high-resolution images with $2048\times 1440$ pixels.
Performance ranking employs the Bjøntegaard delta rate (BD-Rate) metric bjontegaard2001calculation for comprehensive evaluation.
Figure 8: Reconstructions of our proposed MLICv2, recent learned image compression models jiang2023mlicpp, he2022elic, zou2022the, cheng2020learned, feng2025linear, jiang2024llic and VTM-17.0 Intra bross2021vvc. “[MSE]” denotes the model is optimized for MSE. Please zoom in for better view.
(1) In MLICv2+∗, gradient checkpointing chen2016training is employed during the SGA iterations.
(2) The MACs of MLICv2+ are theoretically estimated based on the number of SGA iterations, following the observation from Epoch AI that backward and forward passes almost always have MAC ratios close to 2:1444https://epoch.ai/blog/backward-forward-FLOP-ratio. The MACs of MLICv2+∗ are not reported due to the use of gradient checkpointing, which prevents accurate profiling.
Figure 9: GPU memory consumption, Forward kMACs/pixel, encoding time and decoding time comparisons among our proposed models and recent learned image compression models  cheng2020learned, xie2021enhanced, wang2022neural, zou2022the, he2022elic, liu2023learned, li2024frequencyaware, jiang2022mlic, jiang2023mlicpp.
@@ 4.3.1 Quantitive Results
For a fair comparison with learned image compression models, VTM-17.0 Intra is evaluated in YUV444 color space under encoder_intra_vtm.cfg. To demonstrate the superiority of our MLICv2 and MLICv2+, we compare with recent learned image compression models li2020efficient, cheng2020learned, wu22block, minnen2020channel, qian2020learning, xie2021enhanced, chen2021nic, guo2021causal, qian2022entroformer, zhu2022transformerbased, zou2022the, he2022elic, koyuncu2022contextformer, fu2023learned, feng2023nvtc, liu2023learned, li2024frequencyaware, jiang2024llic, fu2024weconvene, feng2025linear, lu2025learned.
The rate-distortion curves are presented in Fig. 7. When compared with our baseline MLIC++, MLICv2+ performs significantly better, especially at high bit-rates and on high-resolution images tecnick2014TESTIMAGES, CLIC2020. Specifically, our MLICv2 achieves a maximum improvement of $0.6,0.5,0.5$ dB when the bpp is around $0.8,0.55,0.6$ on Kodak, Tecnick, CLIC Pro Val, respectively. Our MLICv2+ achieves larger improvement compared to MLIC++ jiang2023mlicpp due to adopted latent refinement yang2020improving for enhanced instance adaptability. The average improvements of latent refinement on MLICv2 on Kodak kodak, Tecnick tecnick2014TESTIMAGES, CLIC Pro Val CLIC2020 are $0.22,0.15,0.15$ dB, respectively.
BD-Rate reductions over VTM-17.0 Intra are presented in Table 1. Compared with existing methods, our MLICv2 and MLICv2+ achieve state-of-the-art performance, reducing BD-Rates by $16.54\%$ and $20.46\%$ on Kodak. When compared to the baseline model MLIC++ jiang2023mlicpp, our MLICv2 and MLICv2+ reduce $3.15\%$ and $7.07\%$ more bit-rates over VTM-17.0 Intra on Kodak kodak, respectively. Our MLICv2 and MLICv2+ also achieve superior performances on high-resolution datasets Tecnick tecnick2014TESTIMAGES and CLIC Pro Val CLIC2020. Specifically, our MLICv2 and MLICv2+ reduce BD-Rates by $21.61\%$ and $24.35\%$ on Tecnick tecnick2014TESTIMAGES and $16.05\%$ and $19.14\%$ on CLIC Pro Val CLIC2020.
@@ 4.3.2 Qualitative Results
We also compare the qualitative performance to MLIC++, ELIC he2022elic, STF, WACNN zou2022the, Cheng’20 cheng2020learned and VTM-17.0 Intra on subjective qualities. The windowsill of reconstructions are cropped to patches for clearer comparisons. The reconstructions of MLICv2 have sharper textures and retain more details. In terms of visual quality, our MLICv2 has significant improvements on rate-perception performance compared to other models.
@@
4.4 Computational Complexity
The model complexities and sizes are summarized in Fig. 9 and Table 5, respectively. To evaluate complexity across resolutions, 16 images from the LIU4K dataset liu2020comprehensive are cropped to patches of $\{512^{2},768^{2},1024^{2},1536^{2},2048^{2},2560^{2},3072^{2}\}$ pixels, covering a wide range of practical image sizes. We report forward GPU memory, encoding and decoding times (including arithmetic coding), and forward kMACs/pixel as complexity metrics. Since the number of skipped elements varies with bit rate, results for MLICv2 and MLICv2+ are reported at $\lambda=0.0130$.
Due to the high memory consumption of SGA, we adopt gradient checkpointing chen2016training to alleviate GPU memory limitations. Instead of storing all intermediate activations during the forward pass, only a small subset necessary for gradient computation is cached, while the remaining activations are recomputed on-the-fly during backpropagation. This strategy substantially reduces peak memory usage at the cost of increased computation time.
@@ 4.4.1 On Forward GPU Memory
Compared to MLIC+jiang2022mlic, MLICv2 significantly reduces memory usage thanks to its linear-complexity entropy model. Relative to MLIC++jiang2023mlicpp, it incurs slightly higher memory for high-resolution images due to the additional modules introduced in this work, but still requires less memory than Xie’21 xie2021enhanced, Qian’21 qian2020learning, Entroformer qian2020learning, and LIC-TCM liu2023learned. This modest increase is justified by the corresponding performance improvements. Latent refinement, while increasing memory during backpropagation, is optional and primarily benefits scenarios where encoding complexity is less critical. Since images are typically encoded once but decoded multiple times, the added encoding cost is acceptable given the notable gains in rate-distortion performance.
@@
Table 2: Comparisons among different transforms of proposed MLICv2, LIC-TCM liu2023learned, MLIC++, ELIC he2022elic, and Cheng’20 cheng2020learned. All method are quiped with autoregressive context module minnen2018image.
Transform Architecture Cheng’20 cheng2020learned SwinT zhu2022transformerbased ELIC he2022elic MLIC++ LIC-TCM liu2023learned FTIC li2024frequencyaware DCAE lu2025learned MLICv2
BD-Rate(%) 0.00 -4.61 -5.19 -0.41 -8.91 -8.36 -9.17 -9.44
@@
Context Module MLIC++ jiang2023mlicpp MLICv2 HGCP 2D RoPE CR GSC
@@
4.4.2 On Forward kMACs/pixel
Compared to MLIC+, LIC-TCM liu2023learned, our MLICv2 has lower kMACs/pixel. Compared to MLIC++, our MLICv2 exhibits slightly higher kMACs/pixel. This increase can be attributed to two primary factors: the proposed channel reweighting module and selective compression module. However, considering the significant performance gains achieved through these enhancements, the marginal increase in MACs is deemed acceptable. The trade-off between computational complexity and improved compression efficiency is well-balanced in our proposed model.
@@ 4.4.3 On Encoding Time and Decoding Time
Compared to MLIC and MLIC+ jiang2022mlic, MLICv2 achieves higher encoding and decoding efficiency due to its linear-complexity components. Relative to MLIC++jiang2023mlicpp, it further reduces encoding and decoding time by efficiently skipping many zero elements, which is particularly beneficial for high-resolution images where zeros are more frequent. Latent refinement, while involving multiple backpropagation steps and increasing encoding time, is optional and suited for scenarios where encoding complexity is less critical. Since images are typically encoded once but decoded multiple times, the extra encoding cost is justified by the performance gains. Notably, MLICv2+ maintains decoding complexity nearly identical to MLICv2, with both substantially faster than MLIC++, recent WeConvene fu2024weconvene and DCAE lu2025learned.
@@ 4.4.4 On Model Size
Table 4: Complete ablation study on Kodak dataset. BD-Rate (%) is employed to evaluate component contributions. MLICv2+ is the anchor. Cases 1–7 show cumulative benefits, Cases 9–15 show individual contributions. The original Cases 8 and 16 are merged as a separate baseline.
Context Modules MLICv2+ Cumulative Benefits Individual Contributions Baseline
Full Case 1 Case 2 Case 3 Case 4 Case 5 Case 6 Case 7 Case 8 Case 9 Case 10 Case 11 Case 12 Case 13 Case 14 Case 15
ILR
GSC
@@
CR
HGCP
STMT
@@
BD-Rate (%) 0.00 0.17 4.89 5.21 5.77 7.79 8.29 9.58 4.64 9.52 9.12 7.64 8.77 8.51 9.58 9.87